# MultiStream-LLM: Bridging Modalities for Robust Sign Language Translation

### v1

@@ Introduction
Early methods focused on Isolated Sign Language Recognition (ISLR) but relied on expertly annotated datasets. These annotations include “glosses (Camgoz et al. 2018)” which are English translations for individual words. More recent works have demonstrated that glosses are no longer a strict requirement for effective SLT with the introduction of large, gloss-free datasets (e.g. YouTube‐ASL (Uthus et al. 2023), Public DGS (Hanke et al. 2020), How2Sign (Duarte et al. 2021)) as well as increased model size and the use of LLM’s (Lin et al. 2023). However even with ChicagoFSWildPlus (Shi et al. 2019) a dataset focused on ASL fingerspelling modern approaches often struggle with two key challenges:
@@
In parallel, breakthroughs in Automatic Speech Recognition (ASR) (Long et al. 2024) and Visual Speech Recognition (VSR, i.e., lipreading) (Thomas, Fish, and Bowden 2024) demonstrate how large scale pretraining and powerful architectures such as ViTs (Dosovitskiy et al. 2021) and LLM decoders (Radford et al. 2023) can deliver human‐level performance on spoken language tasks. Yet sign language systems lag behind, largely because (a) large, freely available signed corpora are scarce; (b) the spatiotemporal complexity of co‐articulated gestures exceeds that of spoken‐language signals; and (c) unlike speech, signing and mouthing are often offset in time especially during fingerspelling, where manual letters may precede or follow their mouthings.
@@
    Isolated Sign Language Recognition (ISLR): Early work focused on recognizing individual signs in controlled settings. Methods such as 3D CNNs over spatiotemporal volumes (Min, Lee, and Kim 2021) and hybrid CNN–HMM pipelines (Koller, Zargaran, and Ney 2020) achieved high accuracy on small vocabularies (e.g., 100–500 signs), but relied on carefully segmented clips and extensive manual gloss annotations.
@@
    Continuous Sign Language Recognition (CSLR): To scale beyond isolated signs, CSLR methods adopted sequence‐level training with Connectionist Temporal Classification (CTC) loss (Graves et al. 2006a). Recurrent architectures (CNN–BiLSTM–CTC) (Graves and Schmidhuber 2005) and, more recently, Transformer encoders (Min, Lee, and Kim 2021) enabled end‐to‐end gloss prediction from unsegmented video. While these models tolerate variable sign durations, they still depend on gloss‐level transcripts and struggle with out‐of‐vocabulary signs and fingerspelling.
@@
    • Gloss‐free SLT bypasses intermediate glosses, learning a direct video‐to‐text mapping. Recent works leverage large‐scale video–text pretraining (Zhou et al. 2023; Lin et al. 2023; Zhou et al. 2024; Tarrés et al. 2023) and multimodal contrastive objectives (Gueuwou et al. 2024) to improve translation fluency. More recent approaches have focussed on leveraging LLM’s language modelling capabilities to incorporate sign features either from RGB (Uthus, Johnson, and Monz 2023), Pose (Fish and Bowden 2024) , or a fusion of both modalities (Li et al. 2024; Wong, Camgoz, and Bowden 2024). However, these models often underperform on fingerspelling and non‐manual cues, since they attempt to learn all visual subtasks in a single end‐to‐end network.
@@
Lipreading, or visual speech recognition (VSR), can complement sign language understanding by capturing silent speech cues. Modern systems employ deep learning-based architectures. LipNet (Assael et al. 2016) introduced an end-to-end model using spatiotemporal CNNs (LeCun et al. 1998; Carreira and Zisserman 2017) and BiLSTMs (Graves and Schmidhuber 2005), achieving state-of-the-art sentence-level lipreading performance. Transformer-based models, such as Lipformer (Xue et al. 2023), further improved contextual understanding across video frames.
Recent advances leverage self-supervised learning and multimodal fusion to enhance lipreading robustness. AV-HuBERT (Shi et al. 2022) demonstrated the efficacy of large-scale audiovisual pretraining, achieving strong performance under challenging conditions (Thomas, Fish, and Bowden 2024). However, lipreading remains underexplored in the context of fingerspelling and full‐sentence SLT, dispite mouthing cues carrying up to 40% of linguistic information in natural sign production (e.g., lexical disambiguation, grammatical markers) (Aparicio et al. 2017).
Figure 1: An overview of our approach. First we extract facial regions from 16 frames of input video. We apply random pixel masking at 50% probability, add positional embedding, and encode visual features via a vision transformer encoder (ViT-base (Serdyuk, Braga, and Siohan 2022)). We implement temporal downsampling via 1D convolution and then use a Connectionist Temporal Classification (CTC) linear head to predict sequences of phonemes. During training, we also fine-tune an LLM (Vaswani et al. 2017) (Llama (Touvron et al. 2023)) to reconstruct words from phonemes with a text-only dataset (Carnegie Mellon University 2014). During inference, the phonemes from the CTC (Graves et al. 2006b) head are processed via the LLM to reconstruct the predicted text. This can be performed end-to-end or in two stages, depending on available resources.
@@
The rise of Large Language Models (LLMs) has revolutionized sign language understanding, fingerspelling detection, and lipreading by leveraging vast linguistic knowledge and multimodal learning capabilities. In SLT, LLMs have been employed to bridge the gap between sign video sequences and natural language text. Newer approaches (Ruiz and Martinez 2024) applied transformer-based architectures to generate coherent text from sign videos, outperforming traditional gloss-based approaches. GFSLT-VLP (Zhou et al. 2023) further enhanced gloss-free SLT performance.
@@
YouTube-ASL (Uthus, Johnson, and Monz 2023) 37.8 24.1 16.9 12.4 -
SignMusketeers (Gueuwou et al. 2024) 41.5 27.2 19.3 14.3 -
Uni-Sign (Li et al. 2024) 40.2 27.1 19.7 14.9 36.0
@@
YouTube-ASL (Uthus et al. 2023):
@@
FS Recognition (Shi et al. 2018) 41.2%
Iterative Attention (Shi, Jones, and Bowden 2019) 46.7%
Pannattee et al. (Pannattee, Forsyth, and Neumann 2021) 48.0%
Gajurel et al. (Gajurel, Prasad, and Betke 2021) 48.36%
@@
FS Attention (Kabade, Shetty, and Gupta 2023) 57.84%
FSS-Net (Shi, Brighton, and Bowden 2022) 64.4%
CtoML (Li, Cui, and Liu 2023) 54.9%
FS PoseNet (Fayyazsanavi, Shah, and Hu 2024) 71.1%
@@ Results
Table 3: Comparison of Letter Accuracy, Sequence Accuracy and BLEU-4 score when different components are removed from the model. The table shows that removing the LLM drops BLEU-4 down to 7.8 and the Letter Accuracy to 58.9%. When the Lipreading is removed the letter accuracy drops down to 54.3% and the BLEU-4 down to 16.6.

### v2

@@ Introduction
Early methods focused on Isolated Sign Language Recognition (ISLR) but relied on expertly annotated datasets. These annotations include “glosses (Camgoz et al. 2018)” which are English translations for individual words. More recent works have demonstrated that glosses are no longer a strict requirement for effective SLT with the introduction of large, gloss-free datasets (e.g. YouTube‐ASL (Uthus, Tanzer, and Georg 2023), Public DGS (Hanke et al. 2020), How2Sign (Duarte et al. 2021)) as well as increased model size and the use of LLM’s (Lin et al. 2023). However even with ChicagoFSWildPlus (B. Shi and Livescu 2019) a dataset focused on ASL fingerspelling modern approaches often struggle with two key challenges:
@@
In parallel, breakthroughs in Automatic Speech Recognition (ASR) (Long et al. 2025) and Visual Speech Recognition (VSR, i.e., lipreading) (Thomas, Fish, and Bowden 2025) demonstrate how large scale pretraining and powerful architectures such as ViTs (Dosovitskiy et al. 2021) and LLM decoders (Radford et al. 2022) can deliver human‐level performance on spoken language tasks. Yet sign language systems lag behind, largely because (a) large, freely available signed corpora are scarce; (b) the spatiotemporal complexity of co‐articulated gestures exceeds that of spoken‐language signals; and (c) unlike speech, signing and mouthing are often offset in time especially during fingerspelling, where manual letters may precede or follow their mouthings.
@@
    Isolated Sign Language Recognition (ISLR): Early work focused on recognizing individual signs in controlled settings. Methods such as 3D CNNs over spatiotemporal volumes (Min et al. 2021) and hybrid CNN–HMM pipelines (Koller et al. 2019) achieved high accuracy on small vocabularies (e.g., 100–500 signs), but relied on carefully segmented clips and extensive manual gloss annotations.
@@
    Continuous Sign Language Recognition (CSLR): To scale beyond isolated signs, CSLR methods adopted sequence‐level training with Connectionist Temporal Classification (CTC) loss (Graves et al. 2006). Recurrent architectures (CNN–BiLSTM–CTC) (Graves and Schmidhuber 2005) and, more recently, Transformer encoders (Min et al. 2021) enabled end‐to‐end gloss prediction from unsegmented video. While these models tolerate variable sign durations, they still depend on gloss‐level transcripts and struggle with out‐of‐vocabulary signs and fingerspelling.
@@
    • Gloss‐free SLT bypasses intermediate glosses, learning a direct video‐to‐text mapping. Recent works leverage large‐scale video–text pretraining (Zhou et al. 2023; Lin et al. 2023; Zhou et al. 2024; Tarrés et al. 2023) and multimodal contrastive objectives (Gueuwou et al. 2025) to improve translation fluency. More recent approaches have focussed on leveraging LLM’s language modelling capabilities to incorporate sign features either from RGB (uthus2024youtubeasl), Pose (Fish and Bowden 2025) , or a fusion of both modalities (Li et al. 2025; Wong, Camgoz, and Bowden 2024). However, these models often underperform on fingerspelling and non‐manual cues, since they attempt to learn all visual subtasks in a single end‐to‐end network.
@@
Lipreading, or visual speech recognition (VSR), can complement sign language understanding by capturing silent speech cues. Modern systems employ deep learning-based architectures. LipNet (Assael et al. 2016) introduced an end-to-end model using spatiotemporal CNNs (Lecun et al. 1998; Carreira and Zisserman 2018) and BiLSTMs (Graves and Schmidhuber 2005), achieving state-of-the-art sentence-level lipreading performance. Transformer-based models, such as Lipformer (Xue et al. 2023), further improved contextual understanding across video frames.
Recent advances leverage self-supervised learning and multimodal fusion to enhance lipreading robustness. AV-HuBERT (Shi et al. 2022b) demonstrated the efficacy of large-scale audiovisual pretraining, achieving strong performance under challenging conditions (Thomas, Fish, and Bowden 2025). However, lipreading remains underexplored in the context of fingerspelling and full‐sentence SLT, dispite mouthing cues carrying up to 40% of linguistic information in natural sign production (e.g., lexical disambiguation, grammatical markers) (Aparicio et al. 2017).
Figure 1: An overview of our approach. First we extract facial regions from 16 frames of input video. We apply random pixel masking at 50% probability, add positional embedding, and encode visual features via a vision transformer encoder (ViT-base (Dosovitskiy et al. 2021)). We implement temporal downsampling via 1D convolution and then use a Connectionist Temporal Classification (CTC) linear head to predict sequences of phonemes. During training, we also fine-tune an LLM (Vaswani et al. 2023) (Llama (Touvron et al. 2023)) to reconstruct words from phonemes with a text-only dataset (Carnegie Mellon University 2014). During inference, the phonemes from the CTC (Graves et al. 2006) head are processed via the LLM to reconstruct the predicted text. This can be performed end-to-end or in two stages, depending on available resources.
@@
The rise of Large Language Models (LLMs) has revolutionized sign language understanding, fingerspelling detection, and lipreading by leveraging vast linguistic knowledge and multimodal learning capabilities. In SLT, LLMs have been employed to bridge the gap between sign video sequences and natural language text. Newer approaches (Ruiz and Martinez 2025) applied transformer-based architectures to generate coherent text from sign videos, outperforming traditional gloss-based approaches. GFSLT-VLP (Zhou et al. 2023) further enhanced gloss-free SLT performance.
@@
YouTube-ASL (Uthus, Tanzer, and Georg 2023) 37.8 24.1 16.9 12.4 -
SignMusketeers (Gueuwou et al. 2025) 41.5 27.2 19.3 14.3 -
Uni-Sign (Li et al. 2025) 40.2 27.1 19.7 14.9 36.0
@@
YouTube-ASL (Uthus, Tanzer, and Georg 2023):
@@
FS Recognition (Shi et al. 2019b) 41.2%
Iterative Attention (Shi et al. 2019a) 46.7%
Pannattee et al. (Pannattee et al. 2021) 48.0%
Gajurel et al. (Gajurel, Zhong, and Wang 2021) 48.36%
@@
FS Attention (Kabade et al. 2023) 57.84%
FSS-Net (Shi et al. 2022a) 64.4%
CtoML (Li et al. 2023) 54.9%
FS PoseNet (Fayyazsanavi, Nejatishahidin, and Kosecka 2023) 71.1%
@@ Results
Table 3: Comparison of Letter Accuracy and BLEU-4 score when different components are removed from the model. The table shows that removing the LLM drops BLEU-4 down to 7.8 and the Letter Accuracy to 58.9%. When the Lipreading is removed the letter accuracy drops down to 54.3% and the BLEU-4 down to 16.6.